# Sentiment Reasoning for Healthcare

### v3

@@ Abstract
Transparency in AI healthcare decision-making is crucial for building trust among AI and users. Incorporating reasoning capabilities enables Large Language Models (LLMs) to understand emotions in context, handle nuanced language, and infer unstated sentiments. In this work, we introduce a new task - Sentiment Reasoning - for both speech and text modalities, along with our proposed multimodal multitask framework and dataset. Sentiment Reasoning is an auxiliary task in sentiment analysis where the model predicts both the sentiment label and generates the rationale behind it based on the input transcript. Our study conducted on both human transcripts and Automatic Speech Recognition (ASR) transcripts shows that Sentiment Reasoning helps improve model transparency by providing rationale for model prediction with quality semantically comparable to humans while also improving model performance (1% increase in both accuracy and macro-F1) via rationale-augmented fine-tuning. Also, no significant difference in the semantic quality of generated rationales between human and ASR transcripts. All code, data (English-translated and Vietnamese) and models are published online: https://github.com/leduckhai/MultiMed.
[…]
@@
Khai-Nguyen Nguyen∗1,8, Khai Le-Duc∗2,3,8, Bach Phan Tat4, Duy Le5, Long Vo-Dang6, Truong-Son Hy7,8 1College of William and Mary, USA 2University of Toronto, Canada 3University Health Network, Canada 4KU Leuven, Belgium 5Bucknell University, USA 6University of Cincinnati, USA 7 University of Alabama at Birmingham, USA 8 FPT Software AI Center, Vietnam knguyen07@wm.edu, duckhai.le@mail.utoronto.ca, thy@uab.edu
**footnotetext: Equal contribution
@@ 1 Introduction
Figure 1: Visualized pipeline for Sentiment Reasoning. Given an input transcript (either human transcript or ASR transcript), the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE) and its rationale (the reason for this label). It comprises of two tasks: Sentiment Classification and Rationale Generation.
In recent years, speech sentiment analysis has emerged as a significant interdisciplinary field at the intersection of natural language processing (NLP), machine learning, and automatic speech recognition (ASR). This field focuses on the automated detection and interpretation of human emotions and attitudes conveyed through speech, overcoming the limitations of text-based sentiment analysis Murugaiyan and Uyyala (2023) by analyzing the attitude expressed through human voice. The global market for sentiment analysis is projected to expand from an estimated value of US\$4 billion in 2023 to US\$10.1 billion by 2030, exhibiting a compound annual growth rate (CAGR) of 14.2% over the forecast period from 2023 to 2030 Inc (2024).
@@
Sentiment analysis plays a pivotal role across diverse domains, with healthcare being particularly significant. In healthcare customer service, it facilitates real-time evaluation of customer satisfaction, enhancing empathetic and responsive interactions Xia et al. (2009); Na et al. (2012). Moreover, sentiment analysis aids in monitoring the emotional well-being of patients Cambria et al. (2012a), including those with mental health issues such as suicide Pestian et al. (2012). However, these studies only work on text-based sentiment analysis rather than speech-based sentiment analysis.
@@
Despite its potential, speech sentiment analysis presents several technical challenges. First, speech signals are inherently noisy Rajnoha and Pollák (2011) and exhibit significant variability due to differences in accents Turan et al. (2020), speaking styles Shafran and Rose (2003), and recording conditions Yamagishi et al. (2010), complicating the extraction of reliable acoustic features Chen et al. (2017). Second, emotions are subjective Wearne et al. (2019), complex Golan et al. (2006), and multidimensional, making accurate categorization difficult even for humans Kuusikko et al. (2009), thereby necessitating the role of explainable artificial intelligence (AI). Third, given the critical nature of healthcare decisions, where errors can have severe consequences, transparency in AI decision-making is essential to build trust among machines, healthcare professionals, and patients Antoniadi et al. (2021).
To tackle these challenges, we introduce a novel multimodal framework for a novel task: Sentiment Reasoning. Reasoning in AI is crucial for sentiment analysis because it enables deeper understanding beyond surface-level sentiment polarity. By incorporating reasoning capabilities, AI can contextualize emotions within broader discourse, account for nuanced expressions, handle ambiguous language, and infer underlying intent or sentiment that may not be explicitly stated. This enhances the accuracy and reliability of sentiment analysis by considering complex human communication factors that simplistic methods may overlook. Our contributions are as follows:
@@
    We introduce a new task: Sentiment Reasoning for both speech and text modalities, along with our dataset MultiMed-SA
@@
All code, data (English-translated and Vietnamese) and models are published online [^1] https://github.com/leduckhai/MultiMed.
@@ 2.1 Data Collection
The dataset employed for constructing the Sentiment Reasoning dataset was VietMed Le-Duc (2024), the largest and publicly accessible medical ASR dataset in the world. We then annotated sentiment labels (POSITIVE, NEUTRAL, or NEGATIVE) and their corresponding rationales (the reason for this label). This dataset comprises real-world doctor-patient conversations covering a broad spectrum of all available ICD-10 codes (all medical topics), wherein doctors diagnose conditions, explain medical issues, and provide treatment advice.
Details of data annotation pipeline, annotation guidelines, data quality control, and data statistics are shown Section B in the Appendix.
@@ 2.2 Data Statistics
Table 1: Distribution of sentiment labels in the dataset.
Table 1 shows the distribution of sentiment labels in the dataset. In total, the dataset comprises 7878 samples with sentiment labels with rationales. This distribution reflects the dataset’s slight emphasis on neutral content, typical in medical conversations involving detailed explanations and advice.
@@ 3.1 Informal Definition
Previous works Wadhwa et al. (2024); Chen et al. (2024); Hsieh et al. (2023); Ho et al. (2022) have shown that rationale-augmented targets consistently improve the performance of Language Models (LMs). Furthermore, having the option to output the model’s rationale during inference is highly desirable in many scenarios.
@@ 3.2 Formal Definition
Let $x^{T}_{1}:=x_{1},x_{2},...,x_{T}$ be an audio signal of length $T$. Let $C$ be the set of all possible sentiment classes, we should build a speech Sentiment Reasoning model $f$ that both estimates the probability $p(c|x^{T}_{1})$ for each $c\in C$ and generates its rationale sequence $r^{M}_{1}$ of $M$ length.
The decision rule to predict a single sentiment class is:
@@ 3.3 ASR Model
By utilizing Bayes’ Theorem, the probability $p(x)$ can be ignored during maximization since it merely serves as a normalization factor and does not affect the outcome:
$\begin{split}p(w_{1}^{N}|x_{1}^{T})&=\frac{p(x_{1}^{T}|w_{1}^{N})p(w_{1}^{N})}{p(x_{1}^{T})}\\ &\propto p(x_{1}^{T}|w_{1}^{N})p(w_{1}^{N})\end{split}$ (4)
Therefore:
$w^{*}=\operatorname{arg}\max_{w_{1}^{N}}\underbrace{p(x_{1}^{T}|w_{1}^{N})}_{\text{acoustic model}}\cdot\underbrace{p(w_{1}^{N})}_{\text{language model}}$ (5)
@@ 4.2.1 Encoder
The encoder architecture is naturally well-suited for Sentiment Classification, which can be reformulated into the classical classification task. To this end, we directly apply a linear classifier to the output of the encoders. However, encoders can not generate rationales.
We use phoBERT (110M params) Nguyen and Nguyen (2020), a version of RoBERTa Liu et al. (2019) pre-trained on 20GB Vietnamese text, and ViHealthBERT (110M params) Minh et al. (2022), phoBERT trained on 32GB of Vietnamese text in the healthcare domain. For ViHealthBERT, we report the syllable version which achieved better performance than the word version.
@@ 4.2.2 Generative Models
Decoder: We use Vistral-7B-Chat Nguyen et al. (2023) and vmlu-llm [^2] https://huggingface.co/vtrungnhan9/vmlu-llm. Both models have Mistral-7BJiang et al. (2023a) as their backbone. These models were chosen based on their performance on the vmlu benchmark (Vietnamese Multitask Language Understanding) [^3] https://vmlu.ai/leaderboard which compares the performance of Vietnamese Large Language Models (LLMs) on various language understanding tasks.
@@ 4.3 Training with Rationale
Our rationale-augmented training methods are based on, to our knowledge, the current state-of-the-art CoT-distillation approaches for each architecture.
@@
(ii) Post-thinking Chen et al. (2024): For decoder-based models, we augment the training targets by append the human rationale to the label (¡LABEL¿ ¡RATIONALE¿) in a single prompt. Previous works have shown that post-thinking achieved impressive performance Chen et al. (2024); Wadhwa et al. (2024) and compared to pre-thinking where the model first generates its chain-of-thought then provide the label (¡RATIONALE¿ ¡LABEL¿), post-thinking is more stable and token-efficient Chen et al. (2024); Wadhwa et al. (2024) as the model suffers less from hallucination, consistently yields better performance and is more resource efficient as users can already retrieve the target label from the first generated token.
@@ 4.4 Rationale Format
We further study the effects of the format of the rationale on the performance of the generative models. In particular, given the human rationale and human label, we further prompt GPT-3.5-turbo [^4] https://platform.openai.com/docs/models to enhance the rationale into two different format:
@@ 4.5 Evaluation Metrics
For Sentiment Classification task, we employ both accuracy and F1 score, the harmonic mean of precision and recall. It is particularly useful for evaluating models, as it considers both false positives and false negatives, providing a more comprehensive performance metric than accuracy alone.
For Rationale Generation, we employ ROUGE (Recall-Oriented Understudy for Gisting Evaluation) score Lin (2004), which is a set of metrics used to evaluate the quality of generated text by comparing the overlap of n-grams Ney et al. (1994), word sequences, and word pairs between the generated texts and reference texts. Also, we employ BERTScore Zhang et al. , which is an evaluation metric that leverages pre-trained BERT embeddings Devlin et al. (2019) to compute cosine similarity scores between generated texts and reference texts, capturing contextual and semantic nuances more effectively than traditional count-based metrics like ROUGE. BERTscore correlates well with human judgment on sentence-level as proved by Zhang et al. .
@@ 5 Results and Analysis
1. Encoders are efficient yet effective sentiment classification baselines: Encoder models yields the best performance compared to their encoder-decoder and decoder counterparts, with high accuracy scores (¿ 0.665) and stable F1 scores (macro F1 of both models ¿ 0.665). Furthermore, compared to other models, encoders are more parameter efficient which makes them highly economical for training and deployment. We further observe that domain-specific encoders yield notably better performance, with ViHealthBERT outperforming phoBERT in accuracy (+0.8%) and macro F1 (+0.9%).
2. Rationale-augmented training improve model performance: Consistent with previous findings, performing CoT-augmented training on both encoder-decoders and decoders improve our models performance compared to the baseline. Aside from ViT5, other models seem to have benefited from rationale-augmented training, especially the decoder-based LLMs. We observe an average accuracy gain of approximately +0.8% and average macro-F1 gain of approximately +1.3% for these models compared to their label-only counterparts. We further conducted a Student’s t-test Student (1908) and found that the gains are statistically significant for $\alpha=0.1$. This pattern holds for the results in Table 4. We observe a decline in all of our models performance on ASR data which is anticipated due to its WER of 29.6 %. Nonetheless, the models trained with rationale perform noticeably better than models without, with an average accuracy gain of +0.85% and macro F1 gain of +1.4%.
3. The format of post-thinking rationale doesn’t affect the generative models performance: We study the effects of the format of post-thinking rationale on the performance of generative models on Table 4 and observe that it is unclear whether there is a performance gain from more elaborated rationales. This result agrees with previous findings Wadhwa et al. (2024). Particularly, we observe that some models trained on the elaborated rationales have better performance (ViT5 and vmlu-llm). However, for the other models, there is a notable decline in all metrics.
4. Models are likely to misclassify POSITIVE and NEGATIVE transcripts as NEUTRAL: We study the confusion matrix of our best model on human transcript, Vistral7B finetuned with human rationale, on Figure 2. We observe a notable misclassification tendency between NEUTRAL and the other two classes (23.43% and 27.08% with NEGATIVE and POSITIVE respectively). On the other hand, we found that models can easily distinguish NEGATIVE transcripts from POSITIVE ones. This reflects the ambiguity of sentiment analysis data, as in certain cases there is not enough distinctive features that differentiate neutral from other sentiments. Furthermore, given the slightly imbalanced nature of our dataset with fewer POSITIVE examples, its average F1 score is the lowest among the three labels across all models.

### v4

@@ Abstract
Transparency in AI healthcare decision-making is crucial. By incorporating rationales to explain reason for each predicted label, users could understand Large Language Models (LLMs)’s reasoning to make better decision. In this work, we introduce a new task - Sentiment Reasoning - for both speech and text modalities, and our proposed multimodal multitask framework and the world’s largest multimodal sentiment analysis dataset. Sentiment Reasoning is an auxiliary task in sentiment analysis where the model predicts both the sentiment label and generates the rationale behind it based on the input transcript. Our study conducted on both human transcripts and Automatic Speech Recognition (ASR) transcripts shows that Sentiment Reasoning  helps improve model transparency by providing rationale for model prediction with quality semantically comparable to humans while also improving model’s classification performance (+2% increase in both accuracy and macro-F1) via rationale-augmented fine-tuning. Also, no significant difference in the semantic quality of generated rationales between human and ASR transcripts. All code, data (five languages - Vietnamese, English, Chinese, German, and French) and models are published online.
@@
Khai-Nguyen Nguyen∗1, Khai Le-Duc∗2,3, Bach Phan Tat4, Duy Le5, Long Vo-Dang6, Truong-Son Hy7 1College of William and Mary, USA 2University of Toronto, Canada 3University Health Network, Canada 4KU Leuven, Belgium 5Bucknell University, USA 6University of Cincinnati, USA 7University of Alabama at Birmingham, USA [Uncaptioned image]knguyen07@wm.edu       [Uncaptioned image]duckhai.le@mail.utoronto.ca [Uncaptioned image]https://github.com/leduckhai/Sentiment-Reasoning
(*)(*)footnotetext: Equal contribution
@@ 1 Introduction
Figure 1: Visualized pipeline for Sentiment Reasoning  . Given an input transcript (either human transcript or ASR transcript), the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE) and its rationale (the reason for this label). It comprises of two tasks: (1) Sentiment Classification and (2) Rationale Generation . Traditional sentiment analysis only includes Sentiment Classification task, while our framework generates corresponding rationale to explain the reason behind each predicted sentiment label. 9 examples with sentiment labels and their corresponding rationales in our dataset are shown in Table 7 in the Appendix.
@@
Sentiment analysis plays a pivotal role within the healthcare domain. In healthcare customer service, it facilitates real-time evaluation of customer satisfaction, enhancing empathetic and responsive interactions Xia et al. (2009); Na et al. (2012). Moreover, sentiment analysis aids in monitoring the emotional well-being of patients Cambria et al. (2012a), including those with mental health issues such as suicide Pestian et al. (2012). However, these studies only work on text-only sentiment analysis instead of speech-based sentiment analysis.
@@
Despite its potential, speech sentiment analysis presents several technical challenges. First, emotions conveyed through speech are subjective Wearne et al. (2019), complex Golan et al. (2006), and dependent on speaking styles Shafran and Rose (2003), making accurate sentiment classification difficult even for humans Kuusikko et al. (2009), thereby necessitating the role of explainable artificial intelligence (AI). Second, given the critical nature of healthcare decisions, where errors can have severe consequences, transparency in AI decision-making is essential to build trust among machines, healthcare professionals, and patients Antoniadi et al. (2021).
To tackle challenges above, reasoning in AI is crucial for sentiment analysis because it enables deeper understanding beyond surface-level sentiment polarity via the textual explanations. Recent works on Chain-of-Thought (CoT) distillation Wadhwa et al. (2024); Chen et al. (2024); Hsieh et al. (2023); Ho et al. (2022) have revealed that training generative small language models (SLMs) on rationale-augmented targets (the CoT from larger models is provided along side with the target label) can help the SLM (1) perform better and (2) acquire the ability to generate rationale. Our work leverage these findings and prepare a set of human-labeled rationale to train our sentiment analysis models to do Rationale Generation and enhance their performance (Section 4.4 and 4.5). By incorporating rationales to explain reason for each predicted sentiment label, users could understand the model’s reasoning, facilitating better decision-making based on the classification results. Therefore, we introduce a novel multimodal framework for a novel task: Sentiment Reasoning, which comprises of two tasks: (i) Sentiment Classification, in which the model learns to output the sentiment label (POSITIVE, NEUTRAL, or NEGATIVE), and (ii) Rationale Generation, in which the model generates rationale (the free-form text that explains reason for this label). Our contributions are as follows:
@@
    We introduce a new task: Sentiment Reasoning  for both speech and text modalities, along with the world’s largest multimodal sentiment analysis dataset, supporting five languages (Vietnamese, English, Chinese, German, and French)
@@
All code, data and models are published online.
@@ 2.1 Data Collection
The dataset employed for constructing the Sentiment Reasoning  dataset was VietMed Le-Duc (2024), a large and publicly accessible medical ASR dataset. The dataset comprises real-world doctor-patient conversations. We then annotated sentiment labels (POSITIVE, NEUTRAL, or NEGATIVE) and their corresponding rationales (the reason for this label). We then manually translate the transcripts from Vietnamese into other four languages: English, Chinese (Simplified and Traditional), German, and French, making the dataset six times larger. The full dataset (with 5 languages) includes 30000 samples, making it the largest multimodal sentiment analysis dataset, to the best of our knowledge (see Table 2). Our paper focuses mainly on the Vietnamese subset (Section 5) and the English subset (Appendix D).
2.2 Data Annotation
The annotation task consists of two primary steps. First, annotators are required to perform Sentiment Classification. Second, annotators are instructed to provide a rationale behind each class (Rationale Generation ). To ensure consistency, our TESOL-certificated professional linguist has developed an initial guideline inspired by Chen et al. (2020), which was also adopted by various well-known works Shon et al. (2022, 2023), and revised it frequently if necessary. Details of data annotation pipeline, annotation guidelines, data imbalance, translation annotation, and translaton quality control are shown in Appendix Section B.
2.3 Data Quality Control
During the independent annotation process conducted by three annotators, we observed a low inter-annotator agreement (Cohen’s kappa coefficient below 0.5 for the inter-annotator agreement between the two annotators), a common occurrence in real-world datasets as noted by Chen et al. (2020). To address this issue, we implemented an alternative label merging approach. We convened a discussion meeting involving the three annotators and two reviewers (one professional linguist and one with a biomedical background). Each annotator was required to justify their chosen sentiment label and its corresponding rationale. A label and its rationale were selected based on the consensus of all three annotators and two reviewers, rather than a majority vote, as employed in other studies Aziz and Dimililer (2020); Saleena et al. (2018).
@@ 2.4 Data Statistics
Table 1: Distribution of sentiment labels in the dataset for a single language. The real size of the dataset is 6 times larger when accounting all 5 languages - English, Chinese (Simplified and Traditional), German, and French.
Dataset Venue #Samp. #Lang. Domain
Mosi Zadeh et al. (2016) IEEE 3k 1 Vlog
CMU-MOSEI Bagher Zadeh et al. (2018) ACL 23k 1 Various
MELD Poria et al. (2019) ACL 13k 1 TV Series
IEMOCAP Busso et al. (2008) Springer 12k 1 General
SEMAINE McKeown et al. (2012) IEEE 1k 1 Simulation
Sentiment Reasoning (ours) - 30k 5 Medical
Table 2: Data statistics comparison based on the number of samples and languages. Our dataset with 5 languages (Vietnamese, English, Chinese, German and French) includes 30000 samples, making it the largest multimodal sentiment analysis dataset.
Table 1 shows the distribution of sentiment labels in the dataset. This reflects the dataset’s slight emphasis on neutral content, typical in medical conversations involving explanations and advice.
It should be noted that the statistics are reported for a single language, meaning that the real size of the dataset is 6 times larger when accounting all 5 languages.
@@ 3.2 Formal Definition
Let $x^{T}_{1}:=x_{1},x_{2},...,x_{T}$ be an audio signal of length $T$. Let $C$ be the set of all possible sentiment classes, we should build a speech-based Sentiment Reasoning  model $f$ that both estimates the probability $p(c|x^{T}_{1})$ for each $c\in C$ and generates its rationale sequence $r^{M}_{1}$ of $M$ length.
The decision rule to predict a sentiment class is:
@@ 4.1 ASR Model
We employed hybrid ASR setup using wav2vec 2.0 encoder Le-Duc (2024) to transcribe speech to text. The final ASR model has 118M trainable parameters and Word-Error-Rate (WER) of 29.6% on the test set. Details of ASR experiments are shown in Appendix C.1.
4.2 End-to-end Sentiment Classification
We fine-tuned two well-known models, PhoWhisper Le et al. (2024) and Qwen2-Audio Chu et al. (2024), for the end-to-end spoken sentiment analysis task. PhoWhisper is trained large-scale ASR training set consisting of 844 hours of Vietnamese audio, while Qwen2-Audio is trained on more than 500 hours of audio. We use the base version of PhoWhisper with 74M parameters, while Qwen2-Audio has 8.2B parameters.
@@ 4.3.1 Encoder
The encoder architecture is naturally well-suited for Sentiment Classification, which can be reformulated into the classical classification task. To this end, we directly apply a linear classifier to the output of the encoders. However, encoders can not generate rationales. As such, they serve as baselines in our experiments.
We use phoBERT (110M params) Nguyen and Nguyen (2020), RoBERTa Liu et al. (2019) pre-trained on 20GB Vietnamese text, and ViHealthBERT (110M params) Minh et al. (2022), phoBERT trained on 32GB of Vietnamese text in the healthcare domain. For ViHealthBERT, we report the syllable version which achieved better performance than the word version.
@@ 4.3.2 Generative Models
Decoder: We use Vistral-7B-Chat Nguyen et al. (2023) and vmlu-llm [^1] https://huggingface.co/vtrungnhan9/vmlu-llm. Both models have Mistral-7BJiang et al. (2023a) as their backbone. These models were chosen based on their performance on the vmlu benchmark (Vietnamese Multitask Language Understanding) [^2] https://vmlu.ai/leaderboard.
@@ 4.4 Training with Rationale
Previous works Wadhwa et al. (2024); Chen et al. (2024); Hsieh et al. (2023); Ho et al. (2022) have shown that rationale-augmented targets consistently improve the performance of generative language models. Our rationale-augmented training methods are based on, to our knowledge, the current state-of-the-art CoT-distillation approaches for each architecture.
@@
(ii) Post-thinking Chen et al. (2024): For decoder-based models, we augment the training targets by append the human rationale to the label (¡LABEL¿ ¡RATIONALE¿) in a single prompt. Previous works have shown that post-thinking achieved impressive performance Chen et al. (2024); Wadhwa et al. (2024) and compared to pre-thinking where the model first generates its CoT then provide the label (¡RATIONALE¿ ¡LABEL¿), post-thinking is more stable and token-efficient Chen et al. (2024); Wadhwa et al. (2024) as the model suffers less from hallucination, consistently yields better performance and is more resource efficient as users can already retrieve the target label from the first generated token.
@@ 4.5 Rationale Format
While the rationale in our dataset were re-labeled by humans, we are also interested in whether a different and more detailed rationale format would help the models learn better. To this end, we further study the effects of the format of the rationale on the performance of the generative models. In particular, given the human rationale and human label, we further prompt GPT-3.5-turbo to enhance the rationale into two different format:
@@ 4.6 Evaluation Metrics
For Sentiment Classification task, we employ accuracy and class-wise F1 score. For Rationale Generation , we employ ROUGE (Recall-Oriented Understudy for Gisting Evaluation) score Lin (2004). Also, we employ BERTScore Zhang et al. which captures the contextual and semantic nuances. BERTscore has shown to correlate well with human judgment.
@@ 5 Results and Analysis
1. Encoders are efficient yet effective Sentiment Classification baselines: Encoder models yields the best performance compared to their encoder-decoder and decoder counterparts, with high accuracy scores (¿ 0.665) and stable F1 scores (macro F1 of both models ¿ 0.665). We further observe that domain-specific encoders yield notably better performance, with ViHealthBERT outperforming phoBERT in accuracy (+0.8%) and macro F1 (+0.9%).
2. ASR errors have a marginally negative impact on Sentiment Classification performance: For a fair comparison in real-world environments, WERs for human annotators on a standard conversational spontaneous English ASR dataset range from 5% to 15% Stolcke and Droppo (2017) while more challenging real-world ASR datasets are between 17% and 31% Mulholland et al. (2016). Given the complexity of real-world medical conversations, WER of 29.6% by our ASR model is within an acceptable range. Despite the WER of 29.6%, the performance drop in macro F1 scores is small (absolute value of only about 5%).
3. Rationale-augmented training improve model performance: Consistent with previous findings, performing CoT-augmented training on both encoder-decoders and decoders improve our models performance compared to the baseline. We further conducted a Student’s t-test Student (1908) and found that the gains are statistically significant for $\alpha=0.1$. This pattern holds for the results in Table 5. We observe a decline in all of our models performance on ASR data which is anticipated due to its WER of 29.6 %. Nonetheless, the models trained with rationale perform noticeably better than models without, with an average absolute accuracy gain of +0.85%, absolute macro F1 gain of +1.4%, and relative macro F1 gain of +2.5%.
4. The format of post-thinking rationale doesn’t affect the generative models performance: We study the effects of the format of post-thinking rationale on the performance of generative models on Table 5 and observe that it is unclear whether there is a performance gain from more elaborated rationales. This result agrees with previous findings Wadhwa et al. (2024).
5. Models are likely to misclassify POSITIVE and NEGATIVE transcripts as NEUTRAL: We study the confusion matrix of our best model on human transcript, Vistral7B finetuned with human rationale, on Figure 2. We observe a notable misclassification tendency between NEUTRAL and the other two classes (23.43% and 27.08% with NEGATIVE and POSITIVE respectively). On the other hand, we found that models can easily distinguish NEGATIVE transcripts from POSITIVE ones. This reflects the ambiguity of sentiment analysis data. Furthermore, given the slightly imbalanced nature of our dataset with fewer POSITIVE examples, its average F1 score is the lowest among the three labels across all models.
@@
Model Acc. F1 Neg. F1 Neu. F1 Pos. Mac F1
PhoWhisper 0.4651 0.4393 0.5277 0.3328 0.4333
Decoder (Label only)
Qwen2-Audio 0.5815 0.5707 0.6150 0.5208 0.5688
Decoder (Label + Rationale)
Qwen2-Audio 0.5884 0.5875 0.6131 0.5337 0.5781
Table 6: Performance of audio language models
7. Results on end-to-end audio language models We report the results for end-to-end spoken sentiment analysis on PhoWhisper Le et al. (2024) and Qwen2-Audio Chu et al. (2024). Based on the results in Table 6, we make two observations: First, the performance of PhoWhisper is sub-optimal which we attribute to the fact that it was pre-trained for ASR-based tasks. Second, we found that rationale-augmented training can also increase the Sentiment Classification performance for audio language models.